OpenStudy (anonymous):

A square wall measures 10ftx10ft. three windows in the wall measure 3 feet by 4 feet. if you are going to paint the wall, how many squares feet of wall space will be painted?

OpenStudy (anonymous):

Each window = 3*4 = 12 ft^2. The wall = 10*10 = 100 ft^2. (Wall area)-3*(Window area)=your answer.

OpenStudy (anonymous):

and that will be what?

OpenStudy (anonymous):

it has to be 64

OpenStudy (cggurumanjunath):

100-3(12)=100-36=64